OpenStudy (anonymous):

solve this equation using any method and explain the steps 3(w-2)^2-1=6

Parth (parthkohli):

= (3w - 6)^2 - 1 = 6 = (3w - 6) ^ 2 = 7 = 3w^2 + 36 - 216w = 7 = (3w)^2 - 216w= -29 = 9w^2 - 216w = -29 Solve it further

OpenStudy (anonymous):

3(w-2)^2=6+1 (w-2)^2=7/3 (w-2)=+(7/3)^(1/2) or, w=2+(7/3)^(1/2) also (w-2)=-(7/3)^(1/2) or, w=2-(7/3)^(1/2) So there are two solutions

jhonyy9 (jhonyy9):

3(w-2)^2 -1 =6 (w -2)^2 =7/3 w-2=+/- sqrt(7/3) w_1,2=2+/- (sqrt21)/3

jhonyy9 (jhonyy9):

- can you prove it for : 3((2+/- (sqrt21)/3 -2)^2 -1 =6 3((+/- (sqrt21)/3)^2 =7 3( 21/9)=7 21/3=7 7=7

jhonyy9 (jhonyy9):

ok ?